GreenHab Report – January 14th

[category GreenHab-report]

Greenhab Officer:

Hermes Bolivar

Environmental Control:

Heating

Cooling w/ ambient air (3 hrs)

40% Shade Cloth on

80% Shade Cloth on

Average Temperatures:

Low: 16.6°C

High: 31.3°C

Hours of Supplemental Light:

5

Daily Water Usage of Crops:

10 gallons

Water in Blue Tank (gallons):

87.3%, 262 Gallons.

Times of Watering for Crops:

9:12

11:40

15:03

17:04

19:02

Changes to crops:

* There are eight trays ready to plant greens, The greens will be plant tomorrow.

* Some crops show better physiologycal conditions, maybe as a result of the drop of the temperature into the greeenhab.

Narrative:

The max temperature of the greenhab was reduced to 31°C, It’s an important reduction. However, it’s important try of reduce to 25°C or keep below to 30°C for avoid damage in the crops.

Rise the times for watering the crops from 3 to 5 times without change the volume of water, with the aim to icrease the moisture avariable on the crop’s soil along the day.

Watered Crew 201 experiments

Harvest: 7g Cilantro

Support/Supplies Needed:

Operations Report – January 14th

SOL: 2

Name of person filing report: Freddy Castaneda

Non-nominal systems: Suit #9 still not working, Suits #3, 6 and 8 under charging review, a Suit charger is needed because there is one that does not work and no spares are available.

Notes on non-nominal systems: Suit #9 was repaired, the electrical connections were fixed and a broken 5A fuse was identified but it could not be changed because there is no spare fuses, everything but the fuse is operational in the suit at this moment. Suits #3, 6 and 8 are under charging review, pending for diagnostic.

A Suit charger and spare ones are needed.

Generator (hours run): 17hr 37min; Turned on last night (13Jan2019) at 17:40; Turned off this morning (14Jan2019) at 11:17 cloudy day…; Turned on tonight (14Jan2019) at 16:45

Solar SOC – Turned on (13Jan2019) 85%; Turned off (14Jan2019) 100%; Turned on (14Jan2019) 88%

Diesel Reading – Just under 50%

Propane Reading – Under 30%, almost 29%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 90%; 480 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 87%; 260 gallons

Water (loft) – At level marker 10

Static to Loft Pump used – Yes; At 8:45 to refill tank

Water Meter: 01400658

Toilet tank emptied: Yes, a few minutes ago and still feels the smell 😉

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes, used for training EVA´s 1 and 2

Hours: 67.7

Beginning charge: 100%

Ending charge: 90%

Currently charging: Yes

Opportunity rover used: Not used

Hours: 45.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes, used for training EVA’s 1 and 2

Hours: 69.7

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: Nothing to report

HabCar used and why, where? Not used

CrewCar used and why, where? Refueled and returned by the Crew 202 with 172540 miles

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Suit #9 has a broken 5A fuse, still not working, Suits 3,6 and 8 under charging review, one charger is not working; Suit #7 has not been reviewed yet; I´m aware to make a diagnosis.

Summary of Hab operations: Some crew members after a hard review, based on their skills on different engineering areas, have identified that the gas-like smell is probably generated by the metal of the stairs, the welding residues and the oxide, we recommend to give a new layer of paint to the staircase to decrease the smell and protect the metal which is showing some signs of corrosion.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: We were informed this morning by Shannon that yesterday the power was tripped to the power stations because the rover plugs were not completely clean before they were plugged back into power, so the rovers were not charged, she fixed it this morning.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Low batteries on two smoke detectors, spare batteries are needed.

Questions, concerns and requests to Mission Support: Nothing to report

Astronomy Report – January 14th

Crew Astronomer: Santiago Vargas Dominguez

Description: Observing run aimed to complement information in Halpha for M31 data (newly processed image attached).

Some data acquired but observation is yet to be completed.

Observation keep active in the skynet observation list.

Objects Viewed: M31, Heart Nebula

Problems Encountered: none

Sol Summary – January 14th

Sol: 2

Summary Title: Settling in

Author’s name: Oscar Ojeda

Mission Status: Nominal

Sol Activity Summary: We’re getting comfortable in the station, we woke up, shared breakfast and talked about some details on the activity we would have today. Time flew as we had 2 EVAs scheduled, the first group left, made an exploration of the region close to campus, and returned to a hot lunch. Afterwards, the second EVA group left and repeated a similar exploration to group 1, consisting mainly of an on-suit performance assessment, and sample collection. The rest of the day was used to keep the station and rest a little.

Look Ahead Plan: Tomorrow the science begins…! the crew will stay at the hab to begin the work on the research, with some of the samples that were taken today.

Anomalies in work: None

Weather: Mostly cloudy (Cumulus), partially sunny at the afternoon. Temperature similar to yesterday.

Crew Physical Status: In good health, however, we’ve had 3 crewmembers with "more-than-normal" sneezing.

EVA: 2, Space suit performance asessment, and sample collection.

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Journalist Report, EVA Report, Mission Plan, HSO Report.

Support Requested: We don’t have much salt left in the kitchen.

Journalist Report – January 13th

Title: Country Roads, take me home…

Author: Oscar Ivan Ojeda Ramirez

Extract from Commander’s log, Sol 1

We’re finally here. After months of preparation, planning, after rising expectations, we’ve made it. Those alien-like red hills, and geometric shapes of cylinders, hemispheres, and cubes, worthy of a Dali painting, are finally a real place, a place to call home. A whole mission is now ahead of us, a mission full of challenges and rewards, of personal and professional growth, and at the end of it, we expect to come back home to share what we learned, so that others may follow the path.

We venture into the unknown not just to discover it, "it", but most importantly to discover ourselves. Ourselves as individuals, as a crew, as a community, and in our particular case, as a nation, is it the first time a 100% Colombian crew takes the station.

Any doubts I might have had to be on this mission were cleared as soon as we made the turn and the habitat appeared on sight. At that moment it was clear to me that the efforts and sacrifices we are taking are going to be worth it.

And they won’t be few. Amongst others, we have for the better or the worst, the eyes of many people upon us, and we’ll have to perform to the best of our abilities, just to match what they expect from us. And living in a desert, far away from everyone, comes with a toll. But each one of us must not forget why we volunteered to be here, and that being here is the most important thing for our lives right now. We’ve had a history behind with most of the crewmembers and I trust each one of them as a perfect fit for their role within the crew.

Despite its beauty, Mars is definitely not the most welcoming place, it’s cold, it’s dry, and when it’s not dry, it’s muddy. Mars wants for many secrets to stay unveiled, and sometimes will put up a fight on that.

With 14 sols worth of challenges, lessons, laughter, probably tears, but for sure growth, on both a personal and professional point of view, we have to be up for it.

We are now on it together, and as a famous Aristotle quote goes, "the whole is greater than the sum of the parts". It was once a dream, now it is our reality, and soon it will be a memory. It better be a wonderful memory.

Per Aspera Ad Astra.

Greenhab Report – January 13th

Crew 203 Greenhab Report 13-Jan-2019

Greenhab Officer: Hermes Bolivar

Environmental Control:

Heating

Cooling w/ ambient air (7 hrs)

40% Shade Cloth on

80% Shade Cloth on

Average Temperatures:

Low: 15.5°C

High: 36.3°C

Hours of Supplemental Light: 5

Daily Water Usage of Crops: 9 gallons

Water in Blue Tank (gallons): 90.6%, 272 Gallons.

Times of Watering for Crops:

15:02

17:44

20:00

Changes to crops:

* there is evidence of burning of plants due to very high temperature.

* some crops will be discarded for the bad physiology conditions.

Narrative:

It’s necessary to change the temperature of Greenhab and keep among 15 to 25°C and wait for a better condition of the crops during the next days. The period of supplementary light was modified from 18:00 to 23:00. Because it’s turned on during the day. The blue tank was filling with water so the final volume of the tanks is the 272 gallons.

Watered Crew 201 experiments

Harvest:

*3g of Carrot.

*27g of Basil

Support/Supplies Needed: n/a

Operations Report – January 13th

Crew 203 Operations Report 13-JAN-2019

SOL: 1

Name of person filing report: Freddy Castaneda

Non-nominal systems: Suit #9

Notes on non-nominal systems: Still pending for diagnostic.

Generator (hours run): 17hr 15min; Turned on last night (12Jan2019) at 17:29; Turned off this morning (13Jan2019) at 09:25; Turned on tonight (13Jan2019) at 17:40

Solar SOC – Turned on (12Jan2019) 83%; Turned off (13Jan2019) 100%; Turned on (13Jan2019) 85%

Diesel Reading – 55%

Propane Reading – 30%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 95%; 505 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 90%; 270 gallons

Water (loft) – At level marker 11

Static to Loft Pump used – Yes; At 16:50 to refill tank

Water Meter: 01400327

Toilet tank emptied: Yes, it was horrible 😉

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes, used for training Crew 203

Hours: 67.5

Beginning charge: 100%

Ending charge: 72%

Currently charging: Yes

Opportunity rover used: Yes, brakes are still not working correctly but it was used for training Crew 203

Hours: 45.9

Beginning charge: 100%

Ending charge: 80%

Currently charging: Yes

Curiosity rover used: Yes, used for training Crew 203

Hours: 69.3

Beginning charge: 100%

Ending charge: 88%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: Nothing to report

HabCar used and why, where? Used two times to get water in town

CrewCar used and why, where? Yes, used to get Crew 202 to town

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Suit #9 still not working; Suit #7 has not been fixed yet; I´m aware of the situation to make a diagnosis.

Summary of Hab operations: Some crew members have detected a low smell of propane when they go upstairs, we made an inspection with the propane gas detector and nothing was detected.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – January 13th

Crew 203 Operations Report 13-JAN-2019

SOL: 1

Name of person filing report: Freddy Castaneda

Non-nominal systems: Suit #9

Notes on non-nominal systems: Still pending for diagnostic.

Generator (hours run): 16hr 06min; Turned on last night (12Jan2019) at 17:29; Turned off this morning (13Jan2019) at 09:35; Turned on tonight (13Jan2019) at 17:40

Solar SOC – Turned on (12Jan2019) 83%; Turned off (13Jan2019) 100%; Turned on (13Jan2019) 85%

Diesel Reading – 55%

Propane Reading – 30%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 95%; 505 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 90%; 270 gallons

Water (loft) – At level marker 11

Static to Loft Pump used – Yes; At 16:50 to refill tank

Water Meter: 01400327

Toilet tank emptied: Yes, it was horrible 🙁

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes, used for training Crew 203

Hours: 67.5

Beginning charge: 100%

Ending charge: 72%

Currently charging: Yes

Opportunity rover used: Yes, brakes are still not working correctly but it was used for training Crew 203

Hours: 45.9

Beginning charge: 100%

Ending charge: 80%

Currently charging: Yes

Curiosity rover used: Yes, used for training Crew 203

Hours: 69.3

Beginning charge: 100%

Ending charge: 88%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: Nothing to report

HabCar used and why, where? Used to get water in town

CrewCar used and why, where? Yes, used to get Crew 202 to town

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Suit #9 still not working; Suit #7 has not been fixed yet; I´m aware of the situation to make a diagnosis.

Summary of Hab operations: Some crew members have detected a low smell of propane when they go upstairs.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Sol Summary – January 13th

Crew 203 Sol Summary Report 13-Jan-2019

Sol: 1

Summary Title: Country Roads, take me home…

Author’s name: Oscar Ojeda

Mission Status: Nominal, Starting simulation tonight.

Sol Activity Summary: We woke up at 7, along with the departing Purdue crew (Boiler up guys…!). We had our pictures taken, did a small change of command "ceremony", and then departed to a dinosaur walk, guided by director Shannon. We waved goodbye as the crew departed, and then started organizing groceries and rooms.

Today we finished the water runs, and should be ready for the entire length of the mission. We started our training with Atila, being it the first time it’s done in Spanish. We also were able to share breakfast and lunch, which I was able to cook (I love cooking on Mars). Finally, our remote astronomer did an amazing observation of M31, which we will share.

Look Ahead Plan: Tomorrow crew 203 will send the Mission Plan, along with the initial HSO report, and biographies with official pictures. We will start our EVA plans with the first runs and to get to know better our routines within the station, planning to start later on our research projects.

Anomalies in work: None

Weather: Sunny and warm, low to no wind, disperse clouds.

Crew Physical Status: In good health

EVA: none

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Astronomy Report, Journalist Report, EVA Request.

Support Requested: None

[end]